OpenStudy (anonymous):

There are 31 nickles, dimes, and quarters in the drawer with a value of $4.70. How many coins of each type are there if there are twice as many quarters as there are dimes?

OpenStudy (phi):

can you write down any equations?

OpenStudy (anonymous):

Well I thought it might be: Nn + Nd + Nq = 31 5Nn + 10Nd + 25Nq = 470 Nd = 3Nq But when I plug the Nd = 3Nq into the first two and then try to match them up, all I get is fractions.

OpenStudy (radar):

Let x = number of dimes then 2x = number of quarters well then we could further say that since there is a total of 31 coins 31-x-2x = number of nickels

OpenStudy (phi):

radar will show an approach. But your original equations look ok except for the last one. there are twice as many quarters as dimes. that means more quarters than dimes: Nq= 2*Nd (and not Nd= 3Nq)

OpenStudy (radar):

Now for the money. 10x +25(2x) + 5(31-x-2x)=470 10x+50X+155-5x-10x=470 45x=315 x=7 (7 dimes) 2x=14 (quarters) 31-7-14=10 (10 nickels) That sums up to 31 coins.

OpenStudy (radar):

Follow the money dimes 70 cents quarters 3.50 nickels 50 cents Totals to $4.70 The money check, the answers are correct.

OpenStudy (radar):

Your approach was appropriate. However, personally I prefer to keep the number of variables (or unknowns) to a minimum. In this case "one" (x)

OpenStudy (radar):

as phi pointed out you Nd=3Nq was incorrect here is why. The problem there are twice as many quarters as dimes. This would be expressed using your variables as Nq=2Nd

OpenStudy (radar):

Inretrospect, you actually kept the unknown limted to one (N)